Write the equation of the parabola that has the same shape as y = x2 but that has a vertex of (-3, -2)

Answer by dabanfield(803)

The standard vertex form for a parabola is y = a*(x-h)^2 + k where (h,k) is the vertex and "a" determines the shape.
The first parabola can be rewritten as y = 1*(x-0)^2 + 0 in standard vertex form (i.e. a = 1, h=0, and k = 0).
Keeping the shape the same but moving the parabola's vertex to (-3,-2) means changing h to -3 and k to -2 and leaving a=1.
So we have:
y = (x+3)^2 - 2 for the equation of the new parabiola.